8) Write the equation of the perpendicular bisector to the line segment With endpoints (4,-3), (2,-2)
Answer #1

soll let assume in Soll to A (4,-3) B (2,-2) Slop of AB is ma 9-4, 2. is me in с m2 -243 -u m, = -1/2 m, my s -1 2 1 1 2 ma =y = 21 - 6+2 5 1 y=2x - 3.5 L